### v4

@@ Appendix C Additional Experiment Results
C.1 Classification Task
We show more details of our struggle classification task about the top-1 accuracy rate for each of the four splits together with the mean and standard deviation (StdDev) over the accuracy of the four splits. Table 7, 8, and 9 show the additional results for binary classification, four-way classification, and four-way to binary classification, respectively.
@@ C.1 Classification Task
C.2 Regression Task
We provide more experimental results on our struggle level regression task with the evaluation metrics of Mean Squared Error (MSE), Mean Absolute Error (MAE), and the top-1 accuracy rate of binary and four-way classification converted from the predicted struggle level regression scores. These evaluation metrics of various deep models on the three activities for split 1 to 4 are shown in Tables 10, 11, 12, and 13, respectively. The mean and standard deviation (StdDev) for these evaluation metrics for the four splits are shown in Table 14.
C.3 Label Distribution Learning Task
We provide more experimental results about voters’ struggle label distribution learning task with the evaluation metrics of Mean Absolute Error (MAE), Spearman’s Rank Correlation (Spearman’s Rho), and the top-1 accuracy rate of binary and four-way classification by selecting the struggle level classes with the highest frequency in the label distributions and using the voters’ mode label as ground truth. These evaluation metrics of various deep models on the three activities for split 1 to 4 are shown in Tables 15, 16, 17, and 18, respectively. The mean and standard deviation (StdDev) for these evaluation metrics for the four splits are shown in Table 19.
C.4 Further Ablation Details
@@ Appendix D Figures and Tables
Figure 11: Comparison of the distributions for binary struggle labels. The three graphs display the frequency distribution of the binary struggle labels (0 for non-struggle, 1 for struggle) including expert annotation (EA), voters mode (Mode), and voters labels distributions (Voters), for the three activities: Pipes-Struggle, Tent-Struggle, and Tower-Struggle, respectively.
Table 7: Experiment results for binary classification task over all splits.
@@
Table 8: Experiment results for four-way classification task over all splits.
@@
Table 10: Experiment results for regression task—split 1.
@@
Table 11: Experiment results for regression task—split 2.
@@
Table 12: Experiment results for regression task—split 3.
@@
Table 13: Experiment results for regression task—split 4.
@@
Table 14: Experiment results for regression task—results across all four splits including mean values and standard deviation (StdDev).
@@
Table 15: Label distribution learning experiment results on split 1.
@@
Table 16: Label distribution learning experiment results on split 2.
@@
Table 17: Label distribution learning experiment results on split 3.
@@
Table 18: Label distribution learning experiment results on split 4.
@@
Table 19: Label distribution learning experiment results—results across all the four splits including mean values and standard deviation (StdDev).
@@
Only Test Train & Test $\Delta$ Mean Only Test Train & Test $\Delta$ Mean

### v5

@@ B.2 Additional Dataset Statistics
To analyse intra-group voter disagreements, we converted the four-way struggle labels into binary labels and examined the distribution of disagreement levels among voters. The disagreement frequency for each video sample was computed by determining the proportion of voters who selected either struggle or non-struggle, relative to the total number of voters. The smaller of these two proportions was designated as the minority group frequency, representing the level of disagreement among voters.
We show the intra-group disagreement in Figure 13 using three histograms to illustrate the distribution of these minority group frequencies across the three activities. The x-axis represents the frequency of the minority group of voters, ranging from 0 to 0.5, while the y-axis indicates the number of video samples. Additionally, a smoothed KDE curve is overlaid to highlight the density of disagreements within the dataset.
@@ Appendix C Additional Experiment Results
C.1 Struggle Modelling Approach Comparison
We also conducted extensive experiments comparing classification and regression results across different deep models to assess the consistency of this trend (see Table 9 and 10). The results show that for the TSN [68] model, there are three instances where classification achieves higher four-way accuracy (Pipes-Struggle and Tower-Struggle) and binary accuracy (Pipes-Struggle), while regression-to-classification slightly outperforms classification in some cases (e.g., binary accuracy for Tower-Struggle and four-way accuracy for Tent-Struggle). However, classification remains the more stable and effective approach overall. Similarly, for the SlowFast-R50 [16] model, classification consistently outperforms regression-to-classification across all activities in binary accuracy and, in most cases, for four-way accuracy. These findings consistently indicate that models trained for classification achieve superior performance compared to those trained for regression, reinforcing the advantage of direct classification training for struggle detection.
Table 9: Comparison of Classification and Regression-to-Classification Accuracy for TSN [68].
Activity Classification Regression-to-Classification
Binary Accuracy Four-Way Accuracy Binary Accuracy Four-Way Accuracy
Pipes-Struggle 74.54 48.59 72.42 45.02
Tent-Struggle 68.05 38.23 68.27 40.95
Tower-Struggle 82.59 59.49 85.35 58.01
Table 10: Comparison of Classification and Regression-to-Classification Accuracy for SlowFast-R50 [16].
Activity Classification Regression-to-Classification
Binary Accuracy Four-Way Accuracy Binary Accuracy Four-Way Accuracy
Pipes-Struggle 78.01 50.77 76.89 45.28
Tent-Struggle 69.56 39.48 69.06 41.63
Tower-Struggle 88.24 63.92 86.89 56.00
C.2 Struggle Classification
We show more details of our struggle classification modelling approach about the top-1 accuracy rate for each of the four splits together with the mean and standard deviation (StdDev) over the accuracy of the four splits. Table 12, 13, and 14 show the additional results for binary classification, four-way classification, and four-way to binary classification, respectively.
@@ C.2 Struggle Classification
C.3 Struggle VLM Baselines
Table 11: Struggle classification results with VLM baseline. Classification accuracy for binary classification, four-way classification and the four-way to binary classification.
Dataset Pipes-Struggle Tent-Struggle Tower-Struggle
Top-1 Accuracy (%) Top-1 Accuracy (%) Top-1 Accuracy (%)
Random 53.38 28.62 53.38 50.59 27.12 50.59 67.86 40.45 67.86
Majority Class 62.60 39.60 62.60 54.91 35.79 54.91 79.74 57.56 79.74
Video-LLaVA [33] 37.40 14.45 37.40 48.03 24.10 48.03 20.26 12.12 20.26
We used pre-trained Video-LLaVA [33] and evaluated the model’s struggle classification performance on our datasets in a zero-shot setting without further fine-tuning. The evaluation results are shown in Table 11.
We can see that the Video-LLaVA [33] VLM model does not perform very well in both binary and four-way struggle classification. The highest binary classification accuracy is still lower than the random level (48.03 vs 53.38), and the situation remains the same on four-way struggle classification, indicating that the Video-LLaVA [33] model struggles to capture struggle-specific visual cues given text prompts and video input. This is reasonable because the VLM model is not trained or fine-tuned specifically on struggle determination tasks, and the struggle is distinct from many video-understanding tasks on which the VLM model has been trained.
We designed structured input prompts to guide the VLM in identifying struggles. Below is an example of the input prompt used for the Tent-Struggle dataset: “USER: This video shows a person pitching a camping tent in a random 10-second temporal window. Please focus on the hand’s movements and object status and pay attention to the visual signs that may indicate a struggle, such as motor hesitation (e.g., stopping or placement indecision), getting stuck, prolonged actions, non-smooth movements, frequent pauses, repeated attempts, or visible signs of frustration (e.g., hand or head movements). Describe what you observe before deciding if the person is struggling or not. ASSISTANT: ”
To further prompt the VLM to classify struggle and struggle level, we add the following text input: “If no struggle indicators are present, state ‘None observed.’ Then, classify the struggle level on a scale from 1 to 4: (1) definitely non-struggle – smooth, confident movements; (2) slightly non-struggle – minor hesitation but controlled actions; (3) slightly struggle – noticeable pauses, reattempts, or mild frustration; (4) definitely struggle – frequent failed attempts, prolonged hesitation, or strong frustration cues. Provide the output in this format: Observation: [description] Struggle Indicators: [if any] Struggle Level: [1, 2, 3, or 4].”
The text output from the Video-LLaVA [33] VLM model reveals that it suffers from hallucination when handling unseen struggle-related data. In many cases, the Assistant’s responses describe events that are not present in the video. This issue likely arises from the model’s tendency to rely more on textual priors than actual visual input. Furthermore, because struggle-related tasks are not explicitly represented in the VLM’s training data and differ from conventional visual tasks, the model may struggle to extract the necessary temporal features for accurately detecting struggle.
C.4 Struggle Regression
We provide more experimental results on our struggle level regression modelling approach with the evaluation metrics of Mean Squared Error (MSE), Mean Absolute Error (MAE), and the top-1 accuracy rate of binary and four-way classification converted from the predicted struggle level regression scores. These evaluation metrics of various deep models on the three activities for split 1 to 4 are shown in Tables 15, 16, 17, and 18, respectively. The mean and standard deviation (StdDev) for these evaluation metrics for the four splits are shown in Table 19.
C.5 Struggle Label Distribution Learning
We provide more experimental results about voters’ struggle label distribution learning modelling approach with the evaluation metrics of Mean Absolute Error (MAE), Spearman’s Rank Correlation (Spearman’s Rho), and the top-1 accuracy rate of binary and four-way classification by selecting the struggle level classes with the highest frequency in the label distributions and using the voters’ mode label as ground truth. These evaluation metrics of various deep models on the three activities for split 1 to 4 are shown in Tables 20, 21, 22, and 23, respectively. The mean and standard deviation (StdDev) for these evaluation metrics for the four splits are shown in Table 24.
C.6 Multiple Frames and Temporal Ordering
@@ Appendix D Figures and Tables
Figure 12: Comparison of the distributions using binary struggle labels. The three graphs display the frequency distribution of the binary struggle labels (0 for non-struggle, 1 for struggle) including expert annotation (EA), voters mode (Mode), and voters labels distributions (Voters), for the three activities: Pipes-Struggle, Tent-Struggle, and Tower-Struggle, respectively.
Figure 13: Intra-group voter disagreements based on the binary struggle labels. The three graphs illustrate the frequency distribution of disagreements after converting the four-way struggle labels into binary labels. Each graph presents a histogram of video sample counts (y-axis) against the frequency of the minority group of voters (x-axis), with a smoothed kernel density estimate (KDE) curve.
Table 12: Experiment results for struggle binary classification over all splits.
@@
MViTv2 [31] 84.62% 94.44% 82.09% 76.00% 84.29±7.67%
Table 13: Experiment results for struggle four-way classification over all splits.
@@
MViTv2 [31] 81.54% 90.74% 82.09% 76.00% 82.59±6.09%
Table 15: Experiment results for struggle regression—split 1.
@@
MViTv2 [31] 0.82 0.78 86.15% 30.77%
Table 16: Experiment results for struggle regression—split 2.
@@
MViTv2 [31] 0.76 0.73 88.89% 37.04%
Table 17: Experiment results for struggle regression—split 3.
@@
MViTv2 [31] 1.19 0.91 76.12% 35.82%
Table 18: Experiment results for struggle regression—split 4.
@@
MViTv2 [31] 1.05 0.89 74.00% 30.00%
Table 19: Experiment results for struggle regression—results across all four splits including mean values and standard deviation (StdDev).
@@
MViTv2 [31] 0.955±0.200 0.828±0.087 81.29±7.33% 33.41±3.54%
Table 20: Struggle label distribution learning experiment results on split 1.
@@
MViTv2 [31] 0.12 0.5535 76.92% 47.69%
Table 21: Struggle label distribution learning experiment results on split 2.
@@
MViTv2 [31] 0.12 0.6738 87.04% 59.26%
Table 22: Struggle label distribution learning experiment results on split 3.
@@
MViTv2 [31] 0.16 0.3775 73.13% 40.30%
Table 23: Struggle label distribution learning experiment results on split 4.
@@
MViTv2 [31] 0.13 0.5008 84.00% 56.00%
Table 24: Struggle label distribution learning experiment results—results across all the four splits including mean values and standard deviation (StdDev).
@@
Only Test Train & Test
$\Delta$ Mean
Only Test Train & Test
$\Delta$ Mean